OpenStudy (anonymous):

I'm so confused. The graph of a function y = f(x) is shown. Sketch the graph of y = f(x) + 2.

OpenStudy (amistre64):

y = f(4) = 7 then y = f(4) + 2 = 7+2 it just moves the graph up by 2 points everywhere

OpenStudy (anonymous):

The answer is: 

OpenStudy (anonymous):

so you just plot any number for x?

OpenStudy (amistre64):

"The graph of a function y = f(x) is shown" so take that graph and move it up at all points by 2

OpenStudy (amistre64):